# Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations

### v1

@@ Abstract.
Recent advancements in Large Language Models (LLMs) have shown significant potential in enhancing recommender systems. However, addressing the cold-start recommendation problem, where users lack historical data, remains a considerable challenge. In this paper, we introduce KALM4Rec (Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations), a novel framework specifically designed to tackle this problem by requiring only a few input keywords from users in a practical scenario of cold-start user restaurant recommendations. KALM4Rec operates in two main stages: candidates retrieval and LLM-based candidates re-ranking. In the first stage, keyword-driven retrieval models are used to identify potential candidates, addressing LLMs’ limitations in processing extensive tokens and reducing the risk of generating misleading information. In the second stage, we employ LLMs with various prompting strategies, including zero-shot and few-shot techniques, to re-rank these candidates by integrating multiple examples directly into the LLM prompts. Our evaluation, using a Yelp restaurant dataset with user reviews from three English-speaking cities, shows that our proposed framework significantly improves recommendation quality. Specifically, the integration of in-context instructions with LLMs for re-ranking markedly enhances the performance of the cold-start user recommender system.
@@ 1. Introduction
Recommender systems play a pivotal role in assisting users in navigating the vast array of available choices in the digital world. However, a significant and ongoing challenge in this field is addressing the issue of cold-start users. These are users who are new to the platform and therefore have little to no interaction history. The lack of data makes it difficult for the system to generate accurate and personalized recommendations. Conventional Collaborative Filtering (CF) such as BPR (Rendle et al., 2009), NCF (He et al., 2017) and LightGCN (He et al., 2020) or generative methods (Liang et al., 2018; Wang et al., 2023b; Liu et al., 2023) struggle to suggest relevant items effectively for these new users due to the lack of detailed preference information. While user-user content-based algorithms offer a solution by relying on user features to find similar users and recommend positively interacted items, this approach raises privacy concerns. Overall, there remains a significant gap in research addressing the needs of cold-start users, highlighting the necessity for further exploration in this area.
@@
Meanwhile, Large Language Models (LLMs) have garnered considerable interest for their potential to enhance recommender systems, showcasing significant proficiency in natural language understanding, generation, and complex reasoning. Leveraging their robust architecture and extensive training on vast datasets, several studies employ LLMs to comprehend and address user requirements throughout the recommendation process (Gao et al., 2023; Wang et al., 2023a; Li et al., 2024; Wu et al., 2023). Moreover, researchers have explored LLMs’ potential to alleviate the cold-start recommendation problem through in-context learning (He et al., 2023). With LLMs’ adaptability to new tasks, prompts are often tailored to incorporate user and item information, facilitating effective adaptation and performance enhancement (Sanner et al., 2023; Wang et al., 2024).
@@
Approach. For recommender systems on the restaurant domain, user reviews offer valuable insights into their opinions about various businesses. Modeling these personal reviews is beneficial for comprehending users’ preferences and capturing the characteristics of restaurants (Vu et al., 2020). For cold-start users who lack historical data, a practical strategy is to ask them to provide a few keywords describing their preferences. This approach helps create an initial user profile with minimal effort and without compromising their privacy. Furthermore, using keywords instead of entire reviews to prompt LLMs can significantly enhance the efficiency and accuracy of the recommendations while minimizing token usage and computational costs.
To this end, we present a novel framework called Keyword-Driven Retrieval-Augmented Large Language Models for Cold-Start User Recommendations (KALM4REC). Our framework is built upon two essential components centered around keywords: candidates retrieval focusing on retrieving relevant restaurants, and LLM-ranker which leverages LLM to re-rank the retrieved candidates (illustrated in Figure 1). User reviews reflect user preferences but often contain noise, potentially leading to inaccurate representations of user preferences. By exclusively leveraging extracted keywords, we retain the contextual essence of the reviews while still describing the user profile well. Utilizing keywords also allows us to optimize the prompts by reducing noise, thereby enhancing their effectiveness while still ensuring sufficient information of users and restaurants in LLMs, effectively addressing cold-start issues. Essentially, KALM4REC, grounded in sets of keywords, aims to narrow down the candidate set relevant to predefined cold-start preferences. The LLM utilizes its contextual understanding and reasoning capabilities to generate a ranked list of restaurants for recommendations. Our contributions can be summarized as follows:
  • To the best of our knowledge, KALM4REC stands as the pioneering framework to employ keyword as the focal point of both the candidate retrieval and LLM-based re-ranking modules in recommender system. By prioritizing keywords, our augmented LLMs framework effectively addresses the significant challenge of hallucination outcome and optimizing prompt design, resulting in promising performance for cold-start users recommendation.
  • We propose approaches that take advantages of the extracted keywords to obtain most possible restaurants for cold-start users. Furthermore, we designed prompting templates to enable LLMs work efficiently for re-ranking tasks.
  • Finally, we conduct comprehensive experiments to validate and demonstrate the effectiveness of KALM4REC, compared to previous baseline methods for the cold-start user recommendation challenge.
@@ 2. Related works
Cold-start Recommendations. A recommendation system that aims to provide suggestions to new users with little or no historical data is known as a cold-start recommendation system. These systems play a critical role in businesses and services that require recommendations for new users or products. To address the challenge of generating precise recommendations for such cases, various approaches have been proposed. While collaborative filtering (CF) (He et al., 2017; Wang et al., 2019a) has been effective in traditional recommendation systems, it struggles with the cold-start scenario that arises when dealing when new users or items onboard the system. Previous CF-based methods (Bianchi et al., 2017; Gao et al., 2018) rely on data augmentations to incorporate users’ and items’ information into CF to deal with this problem. DropoutNet (Volkovs et al., 2017) employs dropout to directly train the model for cold-start issues, avoiding the need to aggregate all content-based information into objective terms. (Zhao et al., 2022) utilizes latent variables to develop a probabilistic approach for learning a distribution over the side information of items to warm up cold item ID embedding. (Wang et al., 2019b) consider knowledge graph as the source of side information to assist recommendation task. Although these methods show improvements in cold-start problems, they rely heavily on auxiliary information which usually causes privacy concerns.
LLMs for Recommendations. Meanwhile, Large Language Models (LLMs) have attracted significant interest for their ability to improve recommender systems, demonstrating notable skill in natural language understanding, generation, and complex reasoning. Capitalizing on their robust architecture and extensive training on vast datasets, various studies have utilized LLMs to understand and meet user needs throughout the recommendation process (Gao et al., 2023; Wang et al., 2023a; Li et al., 2024; Wu et al., 2023). Additionally, researchers have investigated the potential of LLMs to address the cold-start recommendation problem through in-context learning (He et al., 2023). Given LLMs’ adaptability to new tasks, prompts are often customized to incorporate user and item information, thereby enhancing their effectiveness and performance (Sanner et al., 2023; Wang et al., 2024). Hou et al. (2024) delves into the exploration of LLMs’ capabilities as zero-shot rankers for sequential recommendation. They achieve this by formulating recommendation as a conditional ranking task based on sequential interaction histories, employing meticulously crafted prompt templates, and conducting experiments on some benchmark datasets. Their findings demonstrate the promising zero-shot ranking power of LLMs, surpassing the performance of traditional models. Sanner et al. (2023); Wang et al. (2024) have investigated the inclusion of user and item information in prompts to address cold start challenges. However, none of those previous work have explored the use of keywords to augment LLMs in recommending to cold-start users.
@@ 3. Proposed Framework
In this section, we present the details of our keyword-driven framework, KALM4REC, designed to tackle the cold-start recommendation problem. KALM4REC unfolds in two primary stages: (1) Candidates retrieval, where meaningful noun phrases are extracted from collected reviews to form word sets representing users and restaurants, and subsequently utilized for queries from cold-start users to obtain potential candidates. This phase employs our retrieval models, leveraging a collaborative filtering mechanism with keywords at its core to rank potential candidates; and (2) LLM-based candidates re-ranking, where an LLM is employed as a ranker to re-rank the retrieved candidates.
@@ 3.1. Problem Formulation
Given the set of users $\mathcal{U}$ and the set of restaurants $\mathcal{R}$, the interaction between a user $u\in\mathcal{U}$ and a restaurant $r\in\mathcal{R}$ is presented as a set of reviews =$\mathcal{I}=\{i_{u,r}\}$, where each entry $i_{u,r}$ represents a review if user $u$ has a review for restaurant $r$. The task involves recommending relevant restaurants to a new user $u_{c}$ (cold-start user). The new user can opt to declare his/her preference via a set of pre-defined keywords $k_{u_{c}}$. Simultaneously, user reviews also contain “keywords” $\mathcal{K}=\{k_{u,r}\}$, where each entry $k_{u,r}$ is a set of keywords extracted from $i_{u,r}$. We denote $k_{u}$ as a set of keywords extracted from all user $u$’s reviews, and $k_{r}$ as a set of keywords extracted from all reviews written for restaurant $r$. The problem of keyword-driven cold-start user recommendation is defined as follows. Given a set of keywords $k_{u_{c}}$ from a cold-start (new) user, return a ranked list of relevant restaurants $R_{u_{c}}$ where $R_{u_{c}}\subset\mathcal{R}$. In the following, we now present KALM4REC with the details of the two stages.
@@ 3.2. Candidates Retrieval
In this stage, we present methodologies for retrieving potential candidates that include Content-based Recommendation by Keywords (CBR) and Message Passing on Graph (MPG). For CBR, we utilize deep neural networks to learn the interaction function from sets of keywords representing users and restaurants. To accommodate the arbitrary order of keywords within sets, we adopt a pooling layer based on an attention mechanism. For MPG, we construct a heterogeneous graph comprising nodes representing keywords and restaurants, with two types of edges defined, and introduce a scheme for estimating the scores of the connections between keywords and restaurants.
@@ 3.2.1. Content-based Recommendation by Keywords (CBR)
Motivated by the effectiveness of existing works (Chen et al., 2022), we generate the embedding for the user and restaurants from its keyword representation. For CBR, profile of all users includes cold-start user is presented as sequence of keywords.
Given the sequence of keywords for user $u$: $E_{u}=[E_{u}^{1},E_{u}^{2},...,E_{u}^{T}]\in\mathbb{R}^{T\times d}$, where $T$ denotes the number of keywords for the user and $d$ is the dimension size of the keyword vectorized from sBERT (Reimers and Gurevych, 2019). The recommendation problem can be abstracted as $\hat{y}_{ur}=f(u,r|\Theta)$, where $f$ is a function learned to estimate the score for the user-restaurant pair. Let $p_{u}$ and $q_{r}$ denote the latent vectors for user $u$ and restaurant $r$, learned by $\Theta$. The interaction can be estimated as the inner product of $p_{u}$ and $q_{r}$:
@@
(2) $G(E_{u})=\sigma(W_{c}E_{u}^{T})E_{u}$
where $\sigma$ indicates the softmax layer, and $W_{c}\in\mathbb{R}^{d}$ inferring a trainable parameter. $G(E_{u})$ can be interpreted as a dot product attention, where the keys and the values correspond to the same representation, and the query is only a trainable parameter. Therefore, $G(E_{u})$ represents a weighted average of the encoder sequence of features. The learned features are expected to fully represent the user profile.
For optimization, we adopt Bayesian Personalized Ranking (BPR) (Rendle et al., 2009), where $r_{p}$ is the restaurant that user $u$ prefers over $r_{n}$. The model is optimized by:
@@
where $\hat{y}_{u,r_{p},r_{n}}$ represents the relation between user $u$ and restaurants $r_{p}$ and $r_{n}$; $\sigma$ is the logistic sigmoid function; $\lambda_{\Theta}$ are model regularization parameters.

### v2

@@ Abstract.
Recent advancements in Large Language Models (LLMs) have shown significant potential in enhancing recommender systems. However, addressing the cold-start recommendation problem, where users lack historical data, remains a considerable challenge. In this paper, we introduce KALM4Rec (Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations), a novel framework specifically designed to tackle this problem by requiring only a few input keywords from users in a practical scenario of cold-start user recommendations. KALM4Rec operates in two main stages: candidates retrieval and LLM-based candidates re-ranking. In the first stage, keyword-driven retrieval models are used to identify potential candidates, addressing LLMs’ limitations in processing extensive tokens and reducing the risk of generating misleading information. In the second stage, we employ LLMs with various prompting strategies, including zero-shot and few-shot techniques, to re-rank these candidates by integrating multiple examples directly into the LLM prompts. Our extensive evaluation, using Yelp restaurant data from three English-speaking cities and TripAdvisor hotel data, demonstrates that KALM4Rec excels in improving recommendation quality across the two domains. The framework’s adaptability to different domains highlights its potential for widespread applications. By integrating in-context instructions with LLMs, KALM4Rec notably enhances the performance of cold-start recommender systems, offering a novel approach to exploring solutions in this field. Our code is available at …
@@ 1. Introduction
Recommender systems play a pivotal role in assisting users in navigating the vast number of available choices in the digital world. However, a significant and ongoing challenge in this field is addressing the issue of cold-start users. These are users who are new to the platform and therefore have little to no interaction history. The lack of data makes it difficult for the system to generate accurate and personalized recommendations. Conventional Collaborative Filtering (CF) such as NCF(He et al., 2017), BPR (Rendle et al., 2009)and LightGCN (He et al., 2020) or generative methods (Liang et al., 2018; Wang et al., 2023c; Liu et al., 2023b) struggle to suggest relevant items effectively for these new users due to the lack of detailed preference information. While user-user content-based algorithms offer a solution by relying on user features to find similar users and recommend positively interacted items, this approach raises privacy concerns. Overall, there remains a significant gap in research addressing the needs of cold-start users, highlighting the necessity for further exploration in this area.
@@
Meanwhile, Large Language Models (LLMs) have garnered considerable interest for their potential to enhance recommender systems, showcasing significant proficiency in natural language understanding, generation, and complex reasoning. Leveraging their robust architecture and extensive training on vast datasets, several studies employ LLMs to comprehend and address user requirements throughout the recommendation process (Gao et al., 2023; Wang et al., 2023b; Li et al., 2024; Wu et al., 2023). Moreover, researchers have explored LLMs’ potential to alleviate the cold-start recommendation problem through in-context learning (He et al., 2023). With LLMs’ adaptability to new tasks, prompts are often tailored to incorporate user and item information, facilitating effective adaptation and performance enhancement (Sanner et al., 2023; Wang et al., 2024b).
@@
Approach. For recommender systems in the various domains such as restaurant, user reviews offer valuable insights into their opinions about businesses. Modeling these personal reviews is beneficial for comprehending users’ preferences and capturing the characteristics of restaurants (Vu et al., 2020). For cold-start users who lack historical data, a practical strategy is to ask them to provide a few keywords describing their preferences. This approach helps create an initial user profile with minimal effort and without compromising their privacy. Furthermore, using keywords instead of entire reviews to prompt LLMs can significantly enhance the efficiency and accuracy of the recommendations while minimizing token usage and computational costs.
To this end, we present a novel framework called Keyword-Driven Retrieval-Augmented Large Language Models for Cold-Start User Recommendations (KALM4REC). Our framework is built upon two essential components centered around keywords: candidates retrieval focusing on retrieving relevant items, and LLM-ranker which leverages LLM to re-rank the retrieved candidates (illustrated in Figure 1). User reviews reflect user preferences but often contain noise, potentially leading to inaccurate representations of user preferences. By exclusively leveraging extracted keywords, we retain the contextual essence of the reviews while still describing the user profile well. Utilizing keywords also allows us to optimize the prompts by reducing noise, thereby enhancing their effectiveness while still ensuring sufficient information of users and restaurants in LLMs, effectively addressing cold-start issues. Essentially, KALM4REC, grounded in sets of keywords, aims to narrow down the candidate set relevant to predefined cold-start preferences. The LLM utilizes its contextual understanding and reasoning capabilities to generate a ranked list of items for recommendations. Our contributions can be summarized as follows:
  • To the best of our knowledge, KALM4REC stands as the pioneering framework to employ keyword as the focal point of both the candidate retrieval and LLM-based re-ranking modules in recommender system. By prioritizing keywords, our augmented LLMs framework effectively addresses the significant challenge of hallucinated outcome and optimizing prompt design, resulting in promising performance for cold-start users recommendation.
  • We propose approaches that take advantages of the extracted keywords to obtain most suitable candidates for cold-start users. Furthermore, we designed prompting templates to enable LLMs work efficiently for re-ranking tasks.
  • Finally, we conduct comprehensive experiments to validate and demonstrate the effectiveness of KALM4REC, compared to previous baseline methods for the cold-start user recommendation challenge, using two different domains including restaurant (Yelp) and hotel (Tripadvisor).
@@ 2. Related works
Cold-start Recommendations. A recommendation system that aims to provide suggestions to new users with little or no historical data is known as a cold-start recommendation system. These systems play a critical role in businesses and services that require recommendations for new users or products. Various strategies have been proposed to tackle the challenge of generating accurate recommendations in such scenarios (Guo et al., 2020; Bianchi et al., 2017; Gao et al., 2018; Camacho and Alves-Souza, 2018). Collaborative filtering (CF) (He et al., 2017; Wang et al., 2019a) has traditionally been effective in recommendation systems but struggles with the cold-start problem. To address this challenge, some approaches have integrated data augmentation techniques into CF models, incorporating user and item information to improve recommendations (Bianchi et al., 2017; Gao et al., 2018). While methods such as (Wei et al., 2020; Anwar et al., 2022) extend CF techniques, others like (Son and Kim, 2017) combine collaborative filtering with content-based methods to mitigate the cold-start issue. Deep learning techniques have also been employed to create representations or embeddings that capture the latent features of users and items, aiding in cold-start scenarios (Ying et al., 2018; Volkovs et al., 2017; Zhao et al., 2022; Wang et al., 2019b). DropoutNet (Volkovs et al., 2017) directly addresses cold-start issues by avoiding the need to aggregate all content-based information into the objective function. (Zhao et al., 2022) employs latent variables to develop a probabilistic approach for learning distributions over side information, helping to ”warm up” cold item embeddings. Additionally, graph-based techniques (Ying et al., 2018) leverage user-item interaction graphs to infer user preferences, while (Wang et al., 2019b) uses knowledge graphs as a source of side information to support recommendations. Some methods generate warm embeddings for cold-start items based on their features and ID embeddings, utilizing meta-networks (Zhu et al., 2021). Moreover, (Zhou et al., 2023) explores contrastive learning with CF-based modules to create content-based collaborative embeddings using information from warm-start items to address the cold-start issue. Although these methods show improvements in cold-start problems, they rely heavily on auxiliary information which usually causes privacy concerns.
LLMs for Recommendations.
@@
The impressive capabilities of Large Language Models (LLMs) in natural language understanding, generation, and complex reasoning have inspired innovative approaches in recommendation systems (Fan et al., 2023; Geng et al., 2022; Wang et al., 2024a, 2023b; Li et al., 2024; Wu et al., 2023; Dai et al., 2023; Wang et al., 2023a). Among these, the ”Pretrain, Personalized Prompt, and Predict Paradigm” (P5) (Geng et al., 2022; Xu et al., 2024) stands out as a groundbreaking approach, harnessing the inherent flexibility of the language to tackle diverse recommendation tasks with remarkable success. Rather than fine-tuning models (Lyu et al., 2024; Geng et al., 2022; Cui et al., 2022), some methods use LLMs in a zero-shot setting to recommend items directly, demonstrating promising results (Liu et al., 2023a; Zhang et al., 2023; Kang et al., 2023; Wang and Lim, 2023; Sun et al., 2024). To address the issue of hallucinations when being used for direct recommend ations, LLMs can be employed as re-rankers within recommendation systems(Hou et al., 2024; Dai et al., 2023). For example, Dai et al. (2023) found that GPT could outperform weak baselines when re-ranking a list of candidate items. A notable example of LLMs’ potential is highlighted by Hou et al. (2024), which used LLMs as zero-shot rankers. They transformed the recommendation process into a conditional ranking challenge based on sequential interaction histories and carefully crafted prompt templates. This approach not only demonstrates LLMs’ ability to deliver contextually accurate recommendations but also indicate that by enhancing retrieval modules, LLMs can provide a richer and more responsive user experience. Furthermore, Sanner et al. (2023); Wang et al. (2024b) explore the integration of user and item information into prompts to overcome cold-start issues, while He et al. (2023) shows that in-context learning with LLMs can effectively address these challenges without needing extra data. Despite these advances, there remains a significant gap: no research has yet investigated the combination of keyword extraction with LLMs to boost recommendations for cold-start users. Our study aims to bridge this gap by introducing a novel approach that blends keyword extraction with LLMs to enhance recommendation accuracy in cold-start scenarios.
@@ 3. Proposed Framework
In this section, we present the details of our keyword-driven framework, KALM4REC (Figure 2), designed to tackle the cold-start recommendation problem. KALM4REC unfolds in two primary stages: (1) Candidates retrieval, where meaningful noun phrases are extracted from collected reviews to form word sets representing users and targeted items, and subsequently utilized for queries from cold-start users to obtain potential candidates. This phase employs our retrieval models, leveraging a collaborative filtering mechanism with keywords at its core to rank potential candidates; and (2) LLM-based candidates re-ranking, where an LLM is employed as a ranker to re-rank the retrieved candidates.
@@ 3.1. Problem Formulation
Given the set of users $\mathcal{U}$ and the set of targeted items (items for short) $\mathcal{R}$, the interaction between a user $u\in\mathcal{U}$ and an item $r\in\mathcal{R}$ is presented as a set of reviews =$\mathcal{I}=\{i_{u,r}\}$, where each entry $i_{u,r}$ represents a review if user $u$ has a review for item $r$. The task involves recommending relevant items to a new user $u_{c}$ (cold-start user). The new user can opt to declare their preference via a set of pre-defined keywords $k_{u_{c}}$. Simultaneously, user reviews also contain “keywords” $\mathcal{K}=\{k_{u,r}\}$, where each entry $k_{u,r}$ is a set of keywords extracted from $i_{u,r}$. We denote $k_{u}$ as a set of keywords extracted from all user $u$’s reviews, and $k_{r}$ as a set of keywords extracted from all reviews written for item $r$. The problem of keyword-driven cold-start user recommendation is defined as follows. Given a set of keywords $k_{u_{c}}$ from a cold-start (new) user, return a ranked list of relevant items $R_{u_{c}}$ where $R_{u_{c}}\subset\mathcal{R}$. In the following, we now present KALM4REC with the details of the two stages.
@@ 3.2. Candidates Retrieval
In this stage, we present methodologies for retrieving potential candidates that include Content-based Recommendation by Keywords (CBR) and Message Passing on Graph (MPG). For CBR, we utilize deep neural networks to learn the interaction function from sets of keywords representing users and items. To accommodate the arbitrary order of keywords within sets, we adopt a pooling layer based on an attention mechanism. For MPG, we construct a heterogeneous graph comprising nodes representing keywords and items, with two types of edges defined, and introduce a scheme for estimating scores of the connections between keywords and items.
@@ 3.2.1. Content-based Recommendation by Keywords (CBR)
Motivated by the effectiveness of existing works (Chen et al., 2022), we generate the embedding for the user and items from its keyword representation. For CBR, profiles of all users including cold-start users is presented as sequences of keywords.
Given the sequence of keywords for user $u$: $E_{u}=[E_{u}^{1},E_{u}^{2},...,E_{u}^{T}]\in\mathbb{R}^{T\times d}$, where $T$ denotes the number of keywords for the user and $d$ is the dimension size of the keyword vectorized from sBERT (Reimers and Gurevych, 2019). The recommendation problem can be abstracted as $\hat{y}_{ur}=f(u,r|\Theta)$, where $f$ is a function learned to estimate the score for the user-item pair. Let $p_{u}$ and $q_{r}$ denote the latent vectors for user $u$ and item $r$, learned by $\Theta$. The interaction can be estimated as the inner product of $p_{u}$ and $q_{r}$:
@@
(2) $G(E_{u})=softmax(W_{c}E_{u}^{T})E_{u}$
Here we use a softmax layer, and $W_{c}\in\mathbb{R}^{d}$ inferring a trainable parameter. $G(E_{u})$ can be interpreted as a dot product attention, where the keys and the values correspond to the same representation, and the query is only a trainable parameter. Therefore, $G(E_{u})$ represents a weighted average of the encoder sequence of features. The learned features are expected to fully represent the user profile.
For optimization, we adopt Bayesian Personalized Ranking (BPR) (Rendle et al., 2009), where $r_{p}$ is the item that user $u$ prefers over $r_{n}$. The model is optimized by:
@@
where $\mathcal{D}_{S}=\{(u,r_{p},r_{n})|r_{p}\in\mathcal{I}^{+}_{u}\land j\in\mathcal{I}^{-}_{u}\}$ is the training set, $\mathcal{I}^{+}_{u}$ denoting the items the user $u$ liked and $\mathcal{I}^{-}_{u}$ indicating some items user $u$ doesn’t like; $\hat{y}_{u,r_{p},r_{n}}$ represents the relation between user $u$ and items $r_{p}$ and $r_{n}$; $\sigma$ is the logistic sigmoid function; $\lambda_{\Theta}$ are model regularization parameters.